OpenStudy (anonymous):

Line AB contains points A (8, −4) and B (1, −5). The slope of line AB is

OpenStudy (texaschic101):

To find the slope of line AB, you would have to use the slope formula. Slope = (y2 - y1) / (x2 - x1) (8,-4)...x1 = 8 and y1 = -4 (1,-5)..x2 = 1 and y2 = -5 now we sub slope = (-5 - (-4) / (1 - 8) slope = (-5 + 4) / -7 slope = -1/-7 slope = 1/7 Therefore, the slope of line AB is 1/7

OpenStudy (texaschic101):

Be sure that when you are doing problems like this one, you pay attention to your signs. This is very important.

OpenStudy (anonymous):

yes, k = dy/dx= 1/7 = 0.1429 http://www.hackmath.net/en/calculator/line-slope?x0=8&y0=-4&x1=1&y1=-5&submit=Solve